40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Apr 2 01:18:30 2014, Jenne, Update, LSC, Attempt at PRMI+2 arms 
    Reply  Thu Apr 3 00:30:52 2014, Jenne, Update, LSC, Attempt at PRMI+2 arms 
       Reply  Thu Apr 3 03:06:34 2014, Jenne, Update, LSC, Attempt at PRMI+2 arms 
          Reply  Thu Apr 3 18:09:54 2014, ericq, Update, LSC, Attempt at PRMI+2 arms offsetTRY.pdf2dhist.pdfpdh_trans.pdf
Message ID: 9780     Entry time: Thu Apr 3 00:30:52 2014     In reply to: 9774     Reply to this: 9781
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Attempt at PRMI+2 arms 

[EricQ, Jenne]

We started out this evening by locking the PRMI, on different REFL PDs, just to make sure that we could.  We were able to acquire lock (without having to transition) on either REFL55, REFL33 or REFL165 (I&Q phases for each).  By looking at the transfer function between REFL 55 I and REFL 165 I, I determined that the phase of REFL165 needed to be -135.5, not +44.5, so that the gains were the same sign in all REFL PDs (this is in reference to Koji's elog 9777).  The time to acquisition is much shorter with REFL55 than with the other 2 photodiodes.  I'm not sure right now why this is, but it's pretty consistent, and even more so when the arms are held off resonance.  It is also easy to acquire lock with REFL 55 and then transition to either of the 3f diodes.

PRMI settings:

MICH gain = 2.0.  FMs 4, 5 always on.  FMs 2, 3, 6, 7, 9 triggered with 35 up, 2 down.  Trigger servo on POP110I with 100 up, 10 down.

PRCL gain = -0.020.  FMs 4, 5 always on.  FMs 2, 3, 6, 9 triggered with 35 up, 2 down.  Trigger servo on POP110I with 100 up, 10 down.

PRCL ASC triggered with 50 up, 10 down.  Both pitch and yaw servos had FMs 1, 9 always on, and the FM 6 boost turned on by hand after lock acquired.  Pitch gain = -0.023, yaw gain = -0.027.

If using REFL55, PRCL = I = 1 in the input matrix, and MICH = Q = 1 in the input matrix.

If using REFL33, PRCL = I = 3 in the input matrix, and MICH = Q = 3 in the input matrix.

If using REFL165, PRCL = I = 0.15 in the input matrix, and MICH = Q = 0.1 in the input matrix.


We then moved on to trying to do PRMI + 2 arms, but have been plagued a little bit by locklosses, which may be due to the high seismic from the few large and many small Chilean aftershocks. 

Settings:  With the PRM misaligned, we acquired ALS lock in a CARM/DARM fashion.  The Xarm servo was our "darm" proxy, with +1*ALSX and +1*ALSY.  The Yarm servo was our "carm" proxy, with -1*ALSX and +1*ALSY.  The opposite signs from what you might expect is from our having placed the 2 aux laser frequencies on opposite sides of the PSL.  Our DARM (xarm) was actuating +1 on ETMX and -1 on ETMY.  Our CARM (yarm) was actuating +1 on MC2.  Then we put in a CARM offset of ~60 counts.

We then realigned the PRM, and locked the PRMI (settings as above).  It is much easier to acquire lock with REFL55 than it is with REFL33 or REFL165.  Also, when we are locking the PRMI with REFL55, the AS port is dark, and we see bright sideband resonance at the POP port, as we normally expect.  However, as with last night, when we lock the PRMI with either of the 3f PDs, we see a very bright AS port, and a dark POP port.  Tonight, putting in a MICH offset did not seem to make a significant difference.  As with last night, moving the arms farther away from resonance removed this MICH offset.  I am sure that this is not an artifact of the photodiode dark offsets not being set properly, since I rechecked those carefully. 

We have not come to any interesting conclusions about PRMI+2 arms tonight, since we have started losing the MC every few minutes (maybe seismic-related?). 

ELOG V3.1.3-